# Applying the Viterbi Algorithm to Planetary-Mass Black Hole Searches

### v1

@@ III Test with mock data
Using the python soapcw library soa , we have successfully detected the injected signals that correspond to PBH inspiral signatures. In Fig. 3, we plot the Root Mean Square (RMS), differences between the injected signal and the track detected by the Viterbi algorithm given in units of frequency bins, as functions of the luminosity distance of the source (left panel) and their matched filter SNR (right panel). The larger the RMS value, the larger the deviation between the injected signal and the detected one. As seen in the right panel, we observe that the efficiency of the algorithm is getting progressively worse as we lower the SNR. In the left panel, we observe a similar but reverse correlation between the RMS and the $d_{L}$ of the sources. The result is consistent with our sensitivity estimate, Eq. (32), which indicates $d_{L}=[190,92,36]$kpc (with $n_{\sigma}=10$) for $[10^{-2},5\times 10^{-3},2\times 10^{-3}]M_{\odot}$.
@@
Fig. 4 displays the spectrogram, illustrating the track of the injected signal in Gaussian noise (left panel) and the track detected by Viterbi (right panel) for the case of $[\mathcal{M}c,d_{L}]=[10^{-2}M{\odot},147\mathrm{kpc}]$. The corresponding matched filter SNR is $86.8$. This represents the marginal luminosity distance where the RMS starts to increase for larger distances. We observe that the algorithm successfully traces the signal, with the exception of a small portion at lower frequencies due to the lower SNR in that range. When the source is at a closer distance, the track is accurately detected even at low frequencies.
@@
Furthermore, in Fig. 5, we show a plot of the normalized maximum probability of signal detection via the algorithm as a function of the luminosity distance for the case of $\mathcal{M}_{c}=10^{-2}M_{\odot}$ and $5\times 10^{-3}M_{\odot}$. It is the highest value of the log probability in the last element of the track and can be interpreted as an indicator of confidence. Note that it reflects the extent to which the probability is degraded compared to the noiseless case. Thus, it carries a different meaning from the detection probability. Again, we can see that the algorithms’ efficiency diminishes as the luminosity distance increases. Lastly, in Fig. 6, we present an estimate of the average time required to perform signal detection for $\mathcal{M}c=10^{-2}M{\odot}$. The calculation was conducted on a standard M1 MacBook Pro, utilizing a sample of detections with varying $d_{L}$ values. Note that only the SOAP algorithm part is considered, and the SFT preparation time is excluded. The major advantage of the Viterbi algorithm is evident: the search can be efficiently completed in approximately 10 seconds with a laptop computer. We have observed that the calculation time remains relatively constant across different masses. In fact, the more time-consuming part is the SFT preparation. To prevent signal loss, it is essential to prepare various datasets with differing SFT lengths to search for different PBH masses. Referring to Fig. 2, we find that, to maintain a loss of detection efficiency below $10\%$, the SFTs should be prepared with a mass bin size of $\Delta\log_{10}\mathcal{M}_{c}\sim 0.3$.
@@ IV Conclusions
In the present work, we simulate the data assuming that the noise is stationary and Gaussian. To employ our method in real data analysis, we need to account for the non-stationarity and non-Gaussianity of the detector noise. Particularly, the narrow spectral disturbances, so-called ‘lines’ that can severely degrade the sensitivity Covas et al. (2018). This point remains to be covered in future work. In addition, it would be interesting to cover the cases of asymmetric and extreme mass ratios, as well as to expand the chirp mass range that we consider in our analysis. Currently, the primary bottleneck in terms of computation time is the SFT preparation, as we must adjust the FFT length based on the considered chirp mass. We can tackle this challenge by using SFDBs, incorporating functions that facilitate easy changes to the FFT length. Moreover, BSDs could potentially offer a solution Astone et al. (2005, 2014); Piccinni et al. (2019). This will be another focus of our future work.
@@ Acknowledgements
We acknowledge the i-LINK 2021 grant LINKA20416 by the Spanish National Research Council (CSIC), which initiated and facilitated this collaboration. The authors are grateful to J. Bayley for his useful suggestions regarding the implementation of the SOAP package, and to A. Miller and S. Clesse for very useful discussions. The authors also acknowledge support from the research project PID2021-123012NB-C43 and the Spanish Research Agency (Agencia Estatal de Investigación) through the Grant IFT Centro de Excelencia Severo Ochoa No CEX2020-001007-S, funded by MCIN/AEI/10.13039/501100011033. GA’s and SK’s research is supported by the Spanish Attraccion de Talento contract no. 2019-T1/TIC-13177 granted by the Comunidad de Madrid. GM acknowledges support from the Ministerio de Universidades through Grant No. FPU20/02857. TSY is supported by the JSPS KAKENHI Grant Numbers JP23H04502 and JP23K13099. SK is partly supported by the I+D grant PID2020-118159GA-C42 funded by MCIN/AEI/10.13039/501100011033, the Consolidación Investigadora 2022 grant CNS2022-135211, and Japan Society for the Promotion of Science (JSPS) KAKENHI Grant no. 20H01899, 20H05853, and 23H00110. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.
@@ Appendix A Validity of the leading order post Newtonian expansion
with $t_{c}$ being the coalescence time. From Eq. (39), we can deduce that the the correction to the leading order PN approximation is given by

### v3

@@ III Test with mock data
Using the python soapcw library soa , we have successfully detected the injected signals that correspond to PBH inspiral signatures. In Fig. 6, we plot the Root Mean Square (RMS), differences between the injected signal and the track detected by the Viterbi algorithm given in units of frequency bins, as functions of the luminosity distance of the source (left panel) and their total SNR (right panel). The larger the RMS value, the larger the deviation between the injected signal and the detected one. For a CBC signal, the power deposited in each SFT bin can be very different (the amplitude of the signal drastically increases as the system evolves). Since we expect the Viterbi algorithm to fit better the parts of the track containing more signal, we have defined the RMS differently to Ref. Bayley et al. (2019), weighting by the optimum SNR of each SFT, i.e.
$\mathrm{RMS}=\sqrt{\frac{1}{N_{\mathrm{SFT}}}\sum_{i=1}^{N_{\mathrm{SFT}}}\frac{(\rho_{i}^{\mathrm{opt}})^{2}}{(\rho_{\mathrm{tot}}^{\mathrm{opt}})^{2}}\left(\frac{\nu_{0,i}-\nu_{V,i}}{\Delta f}\right)^{2}}\,,$ (56)
where $\nu_{0,i}$ and $\nu_{V,i}$ denote the frequencies the tracks go through in the $i$-th SFT for the injected signal and the Viterbi track respectively, while $(\rho_{i}^{\mathrm{opt}})^{2}$ denotes the squared optimum SNR in the $i$-th SFT such that $(\rho_{\mathrm{tot}}^{\mathrm{opt}})^{2}=\sum_{i=1}^{N_{\mathrm{SFT}}}(\rho_{i}^{\mathrm{opt}})^{2}$.
As seen in the right panel, we observe that the efficiency of the algorithm is getting progressively worse as we lower the SNR. In the left panel, we observe a similar but reverse correlation between the RMS and the $d_{L}$ of the sources. The result is consistent with our sensitivity estimate, Eq. (32), which indicates $d_{L}=[190,92,36]$kpc (with $n_{\sigma}=10$) for $[10^{-2},5\times 10^{-3},2\times 10^{-3}]M_{\odot}$.
@@
Fig. 7 displays the spectrogram, illustrating the track of the injected signal in Gaussian noise (left panel) and the track detected by Viterbi (right panel) for the case of $[\mathcal{M}c,d_{L}]=[10^{-2}M_{\odot},147\mathrm{kpc}]$. The corresponding matched filter SNR is $86.8$. This represents the marginal luminosity distance where the RMS starts to increase for larger distances. We observe that the algorithm successfully traces the signal, with the exception of a small portion at lower frequencies due to the lower SNR in that range. If the source is at a closer distance, the track is accurately detected even at low frequencies.
Figure 8: For the different chirp mass cases, we show as a function of luminosity distance $d_{L}$ the number of standard deviations $n_{\sigma}$ the SNR is away from the zero signal mean given by Eq. (53) (dots) and approximated expected value of $n_{\sigma}$ of Eq. (54) (solid lines).
In Fig. 8 we show the number of standard deviations $n_{\sigma}$, which indicates the extent to which the observed value deviates from the zero signal mean as given in Eq.(53), as a function of the luminosity distance $d_{L}$ for different chirp masses. The realizations are the same as those in Fig.6. We observe that the approximation in Eq.(54) for the expected value of $n_{\sigma}$ fits the data well at small values of the luminosity distance, i.e., large SNRs. However, as expected, when $(\rho_{\mathrm{tot}}^{\mathrm{opt}})^{2}\to\mu(\rho^{2}_{\mathrm{tot,max,0}})-2N_{\mathrm{SFT}}$, the approximation becomes worse and tends to underestimate $n_{\sigma}$. Comparing with Fig. 6, we observe that the approximation begins to fail when the RMS becomes large. This implies that the Viterbi algorithm is not following the injected track but rather is tracking a noise-dominated path, and the assumption of following the injected track, used to derive our approximation is no longer valid. Looking at the distance at which the data points cross $n_{\sigma}=10$, we observe that the sight distance derived in Eq. (32) provides a reliable estimate of the search range. This indicates that, in terms of $n_{\sigma}$, the Viterbi algorithm does not significantly lose sensitivity due to the trial factor. We have to keep in mind, however, that having the same $n_{\sigma}$ does not mean having the same statistical significance; this depends on the underlying distribution. The distribution of $\rho^{2}_{\mathrm{tot,max,0}}$, shown in Fig. 5, has heavier tails than a Gaussian, and therefore the same value of $n_{\sigma}$ than in Eq. (32) is less significant.
Lastly, in Tab. 1, we present the average time required to perform signal detection for each case of $\mathcal{M}_{c}$ along with the corresponding standard deviation. The calculation was conducted on a standard M1 MacBook Pro, utilizing a sample of detections with varying $\mathcal{M}_{c}$ values. Note that only the SOAP algorithm part is considered here, and the SFT preparation time is excluded.
$\mathcal{M}_{c}(M_{\odot})$ $N_{\mathrm{SFT}}$ $n_{f}$ $\Delta t(s)$
$10^{-2}$ 2560 552 1.34 $\pm$ 0.03
$5\times 10^{-3}$ 2730 1035 1.79 $\pm$ 0.02
$2\times 10^{-3}$ 1280 2208 3.00 $\pm$ 0.01
Table 1: The average time ($\Delta t$) needed in order to complete a signal detection using a standard M1 Macbook Pro for each case of $\mathcal{M}_{c}$ used in our analysis, along with its standard deviation. For reference, we also show the number of SFTs $N_{\mathrm{SFT}}$ and frequencies $n_{f}$ for each case.
@@
The major advantage of the Viterbi algorithm is evident: the search can be efficiently completed in approximately 10 seconds with a laptop computer. In fact, the more time-consuming part is the SFT preparation. To prevent signal loss, it is essential to prepare various datasets with differing SFT lengths to search for different PBH masses. Referring to Fig. 2, we find that, to maintain a loss of detection efficiency below $10\%$, the SFTs should be prepared with a mass bin size of $\Delta\log_{10}\mathcal{M}_{c}\sim 0.3$.
@@ IV Conclusions
In the present work, we simulate the data assuming that the noise is stationary and Gaussian. To employ our method in real data analysis, we need to account for the non-stationarity and non-Gaussianity of the detector noise. Particularly, the narrow spectral disturbances, so-called ‘lines’ that can severely degrade the sensitivity Covas et al. (2018). This point remains to be covered in future work. In addition, it would be interesting to cover the cases of asymmetric and extreme mass ratios, as well as to expand the chirp mass range that we consider in our analysis. Currently, the primary bottleneck in terms of computation time is the SFT preparation, as we must adjust the FFT length based on the considered chirp mass. We can tackle this challenge by using short FFT databases (SFDBs) Astone et al. (2005, 2014), incorporating functions that facilitate easy changes to the FFT length. Moreover, Band Sampled Data (BSD) could potentially offer a solution Piccinni et al. (2019). This will be another focus of our future work.
Numerical Analysis Files: The codes used by the authors in the analysis of the paper will be made publicly available upon publication here.
@@ Acknowledgements
We would like to thank Andrew Miller for his work reviewing this paper within the LIGO and Virgo Collaborations. We acknowledge the i-LINK 2021 grant LINKA20416 by the Spanish National Research Council (CSIC), which initiated and facilitated this collaboration. The authors are grateful to J. Bayley for his useful suggestions regarding the implementation of the SOAP package, and to S. Clesse for very useful discussions. The authors also acknowledge support from the research project PID2021-123012NB-C43 and the Spanish Research Agency (Agencia Estatal de Investigación) through the Grant IFT Centro de Excelencia Severo Ochoa No CEX2020-001007-S, funded by MCIN/AEI/10.13039/501100011033. GA’s and SK’s research is supported by the Spanish Attraccion de Talento contract no. 2019-T1/TIC-13177 granted by the Comunidad de Madrid. GM acknowledges support from the Ministerio de Universidades through Grant No. FPU20/02857. TSY is supported by the JSPS KAKENHI Grant Numbers JP23H04502 and JP23K13099. SK is partly supported by the I+D grant PID2020-118159GA-C42 funded by MCIN/AEI/10.13039/501100011033, the Consolidación Investigadora 2022 grant CNS2022-135211, and Japan Society for the Promotion of Science (JSPS) KAKENHI Grant no. 20H01899, 20H05853, and 23H00110. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.
@@ Appendix A Validity of the leading order post Newtonian expansion
with $t_{c}$ being the coalescence time. From Eq. (57), we can deduce that the correction to the leading order PN approximation is given by
@@
Appendix B Distribution of Viterbi total SNR for uncorrelated paths
Figure 9: Mean $\mu$ and standard deviation $\sigma$ of the total SNR of the Viterbi track as a function of $n_{J}$, computed to leading order in $N_{\mathrm{SFT}}$. We have computed it using Eq. (73)
Here, we provide an analytical estimation of the distribution of the maximum total SNR squared in the case where tracks are not correlated. The cumulative distribution function (CDF) of the maximum $\rho_{\mathrm{tot},\vec{\nu}}^{2}$ of $N_{t}$ independent paths can be computed as
$P_{V}(x)$ $\equiv P\left(\bigcap_{\vec{\nu}}\rho_{\mathrm{tot},\vec{\nu}}^{2}<x\right)=\prod_{\vec{\nu}}P\left(\rho_{\mathrm{tot},\vec{\nu}}^{2}<x\right)$
$=P\left(\rho_{\mathrm{tot}}^{2}<x\right)^{N_{t}}\,.$ (63)
Limiting ourselves to the case with no signal, then $P\left(\rho_{\mathrm{tot}}^{2}<x\right)$ is the CDF of a $\chi^{2}$ distribution with $2N_{\mathrm{SFT}}$ digrees of freedom,
$P\left(\rho_{\mathrm{tot}}^{2}<x\right)=1-\frac{\Gamma(N_{\mathrm{SFT}},x/2)}{\Gamma(N_{\mathrm{SFT}})}\,,$ (64)
where $\Gamma(N_{\mathrm{SFT}},x/2)$ is the upper incomplete gamma function Abramowitz and Stegun (1974). Then we obtain
$P_{V}(x)=\exp\left\{N_{t}\log\left(1-\frac{\Gamma(N_{\mathrm{SFT}},x/2)}{\Gamma(N_{\mathrm{SFT}})}\right)\right\}\,.$ (65)
Now, our interest lies in the center of the probability distribution rather than its tail. Therefore, when $N_{t}\gg 1$, the logarithmic component is a negative number which should be very small. In this regime, we can employ the approximation $\log(1-X)\approx-X$ and proceed to obtain
$P_{V}(x)$ $\approx\exp\left\{-N_{t}\frac{\Gamma(N_{\mathrm{SFT}},x/2)}{\Gamma(N_{\mathrm{SFT}})}\right\}$
$=\exp\left\{-e^{-\alpha(x)}\right\}\,,$ (66)
where we have defined
$\alpha(x)=\log\frac{\Gamma(N_{\mathrm{SFT}})}{N_{t}}-\log{\Gamma(N_{\mathrm{SFT}},x/2)}\,.$ (67)
In the limit of very large $N_{\mathrm{SFT}}$, $\alpha(x)$ can be approximated as Abramowitz and Stegun (1974)
$\alpha(x)$ $\approx\log\frac{\Gamma(N_{\mathrm{SFT}})}{N_{t}}+\frac{x}{2}-(N_{\mathrm{SFT}}-1)\log\frac{x}{2}+O\left(\frac{x}{N_{\mathrm{SFT}}}\right)\,.$ (68)
This $\alpha(x)$ becomes zero when
$x_{0}=2(N_{\mathrm{SFT}}-1)z_{0}\,,\;\mathrm{where}\;\;\log(z_{0})-z_{0}+C=0\,,$ (69)
with $C$ given by
$C$ $=\log(N_{\mathrm{SFT}}-1)\frac{\log(N_{t})-\log(\Gamma(N_{\mathrm{SFT}}))}{N_{\mathrm{SFT}}-1}$
$\approx 1+\log(2n_{J}+1)+O\left(\frac{1}{N_{\mathrm{SFT}}}\right)\,.$ (70)
Here, we have assumed $N_{t}=n_{f}(2n_{J}+1)^{N_{\mathrm{SFT}}}$ with $N_{\mathrm{SFT}}\gg 1$ and $N_{\mathrm{SFT}}\gg\ln(n_{f})$. Therefore, we have $x_{0}\sim O(N_{\mathrm{SFT}})$, and from Eqs. (68) and (69), we obtain
$\alpha^{\prime}(x_{0})\approx\frac{1}{2}\left(1-\frac{1}{z_{0}}\right)+O\left(\frac{1}{N_{\mathrm{SFT}}}\right)\,,$ (71)
and $\alpha^{\prime\prime}(x_{0})\sim O(1/N_{\mathrm{SFT}})\to 0$. Therefore, the CDF of Eq. (66) follows the Gumbel distribution
$P_{V}(x)\approx\exp\left\{-e^{-\alpha^{\prime}(x_{0})(x-x_{0})}\right\}\,,$ (72)
in accordance with the extreme value theorem. This is a well known distribution, having a mean $\mu$ and standard deviation $\sigma$ given by
$\mu=x_{0}+\frac{\gamma}{\alpha^{\prime}(x_{0})}\quad\mathrm{and,}\quad\sigma=\frac{\pi}{\alpha^{\prime}(x_{0})\sqrt{6}}\,,$ (73)
where $\gamma$ is the Euler-Mascheroni constant. In Fig. 9, we show these quantities, to leading order in $N_{\mathrm{SFT}}$, as a function of $N_{\mathrm{SFT}}$. For a value of $n_{J}=1$, which is the number of jumps we allow for in our Viterbi implementation, from Eq. (69), we have $z_{0}(n_{J}=1)=3.289$, and therefore $\mu_{n_{J}=1}\sim 6.58N_{\mathrm{SFT}}$ and $\sigma_{n_{J}=1}=3.69$.